write an equation for the line passing through the point (4,5) and perpendicular to the line passing through the points (-1,3) and (2,9)
Answer by mananth(16946)

x1 y1 x2 y2
-1 3 2 9

slope m = (y2-y1)/(x2-x1)
( 9 - 3 )/( 2 - -1 )
( 6 / 3 )
m= 2
The slope of a line perpendicular to the above line will be the negative reciprocal
m1*m2=-1
The slope of the required line will be - 1/2

m= - 1/2 ,point ( 4 , 5 )
Find b by plugging the values of m & the point in
y=mx+b
5 = -2.00 + b
b= 7
m= - 1/2
The required equation is y=- 1/ 2x+7
m.ananth@hotmail.ca